Determine the maximal domains of the functions f(x) = 1/x−5
and g(x) =
√8-2x

Answer by ikleyn(52946)

.
For the function  f(x) = (1/x) - 5 the domain is the entire number line, except x= 0 where the denominator is zero.


For the function  f(x) = 1/(x - 5) the domain is the entire number line, except x= 5 where the denominator is zero.


For the function  f(x) = sqrt(8-2x)  the domain is the set  8-2x >= 0,   or, which is the same,

8 >= 2x,   or   x <= 4,

providing that the binomial under the square root is non-negative.


------------
Unfortunately, from your writing I can not distinguish which of the two fractions you are asking about,  (1/x) - 5  or  1/(x - 5).

In the future,  when you will post formulas to this forum,  use parentheses to make your formulas unambiguous.

Using parentheses in formulas is  THE RULE  and  THE REQUIREMENT  in this forum.

Thank you.